# SageAttention2: Efficient Attention with Thorough Outlier Smoothing and Per-thread INT4 Quantization

### v3

@@ 3.2 Smooth Q
3.3 INT4 Per-thread Quantization
SageAttention uses per-block quantization, which quantizes each block of $Q$ and $K$ for a GPU streaming processor. Such a quantization strategy could achieve an accuracy performance close to per-token quantization and avoid the overhead of the dot product of the quantization scale vectors $\delta_{Q}$ and $\delta_{K}$. However, quantizing $Q$ and $K$ to INT4 demands a more accurate quantization granularity. We propose per-thread quantization, a more precise and granular quantization approach than the per-block quantizer, also without the additional overhead of vectors dot product required for per-token quantization.
Specifically, each block of $Q$, i.e., $Q_{i}$, in SageAttention will be split into $c_{w}$ segments and processed by $c_{w}$ GPU warps in a GPU streaming processor (SM). We call each segment of $Q_{i}$ as $Q_{w}$, and $k_{w}=K_{j}$ because a $K_{j}$ only corresponds to one warp. Then, each warp containing 32 threads uses the mma.m16n8k64 PTX instruction (NVIDIA, ) for the $Q_{w}K_{j}^{\top}$. According to the layout requirement of this instruction, we find that the $Q_{w}[8\times(n\%8)]$ could share one quantization scale, and $K_{j}[8\times(n\%8)]$ and $K_{j}[8\times(n\%8+1)]$ could share one quantization scale, where $n$ is the token index. Such a quantization method is more fine-grained and with no additional overhead. This is because it assigns different GPU threads to distinct quantization groups based on the MMA instruction layout, with each thread performing dequantization only using a single quantization scale value. The per-thread quantization process can be formulated as follows.
$i_{\delta q}=$ $\lfloor(n*8*c_{w}/b_{q})\rfloor$
$q_{i}[i_{\delta q}]=$$\{8\times(n\%8)+\lfloor(n*c_{w}/b_{q})\rfloor*b_{q}/c_{w}\},~{}~{}n\in[0,N]$
$\hat{Q}[q_{i}[i_{\delta q}]]=$ $\left\lceil\frac{Q[q_{i}[i_{\delta q}]]}{\delta_{Q}[i_{\delta q}]}\right\rfloor$ (5)
$i_{\delta k}=$ $\lfloor(n*4/b_{k})\rfloor$
$k_{n}[i_{\delta k}]=$$\{8\times(n\%8)+\lfloor n/b_{k}\rfloor*b_{k}\}\cup\{8\times(n\%8)+1+\lfloor n/b_{k}\rfloor*b_{k}\},~{}~{}n\in[0,N]$
$\hat{K}[k_{n}[i_{\delta k}]]=$ $\left\lceil\frac{K[k_{n}[i_{\delta k}]]}{\delta_{K}[i_{\delta k}]}\right\rfloor$ (6)
We show an example of per-thread quantization in Figure 6.
Algorithm 1 Implementation of SageAttention2.
  Preprocessing: $K=K-\mathrm{mean}(K)$, $\overrightarrow{Q}_{m}=\mathrm{mean}(Q)$, $Q=Q-\overrightarrow{Q}_{m}$, $\Delta S=\text{GEMV}(\overrightarrow{Q}_{m},K^{\top})$.
  Quantization: ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{Q},\hat{Q})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{Q}}}(Q)$, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{K},\hat{K})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{K}}}(K)$. //per-thread. ${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{V}}}(V)$. //per-channel.
  Divide $\hat{Q}$ to $T_{m}={N}/{b_{q}}$ blocks $\{\hat{Q}_{i}\}$; divide $\hat{K}$, $\hat{V}$, and $\Delta S$ to $T_{n}={N}/{b_{kv}}$ blocks $\{\hat{K}_{i}\}$, $\{\hat{V}_{i}\}$, and $\{\Delta S_{i}\}$;
  for $i=1$ to $T_{m}$ do
     Load $\hat{Q}_{i}$ and $\delta_{Q}=\delta_{Q}[c_{w}*8i:c_{w}*8(i+1)]$ into an SM ;
     for j in [1, $T_{n}$do
        Load $\hat{K}_{j}$, $\hat{V}_{j}$, and $\delta_{K}=\delta_{K}[4j:4(j+1)]$ into the SM ;
        $w=\text{range}(c_{w}),st=w*c_{w}$
        $S_{i}^{j}[st:st+c_{w}]=\psi^{-1}_{\delta_{Q}\delta_{K}}(\mathrm{Matmul}(\hat{Q}_{i}[st:st+c_{w}],\hat{K}_{j}^{\top}))+\Delta S_{j}$; // Paralleled by $c_{w}$ warps. The dequantization $\psi^{-1}_{\delta_{Q}\delta_{K}}$ is illustrated in Figure 6.
        $m_{i}^{j}=\mathrm{max}(m_{i}^{j-1},\mathrm{rowmax}(S_{i}^{j}))$, ${}\widetilde{P}_{i}^{j}=\mathrm{exp}(S_{i}^{j}-m_{i}^{j})$, ${}l_{i}^{j}=e^{m_{i}^{j-1}-m_{i}^{j}}+\mathrm{rowsum}(\widetilde{P}_{i}^{j})$ ;
        $O_{i}^{j}$(FP22) = $\mathrm{Matmul}((\widetilde{P}_{i}^{j}*448)\text{.to(FP8.e4m3)},V_{j})$ ;
        $O_{i}^{j}$(FP32) = $\mathrm{diag}(e^{m_{i}^{j-1}-m_{i}^{j}})^{-1}O_{i}^{j-1}(FP32)+O_{i}^{j}$(FP22) ;
     end for
     Load $\delta_{V}$ into an SM;
[…]
     Write $O_{i}$ ;
  end for
  $O=\{O_{i}\}$
  return $O$
Table 4: Error of the FP8 Matmul instruction of mma(f8f8f32).
Precision of Accumulated Value E8M13 E8M23
Error compared to FP32 0 ||||mma(f32.f16.f16.f32) - mma(f32.f8.f8.f32)||||
3.4 FP32 MMA Buffer for FP22 Accumulator
To avoid the drawback of SageAttention where FP16 Matmul with an FP16 accumulator for $PV$ is only effective on GPUs such as the RTX4090, we adopt the approach of quantizing $P$ and $V$ to FP8 to leverage the universal acceleration of FP8 Tensor cores. However, we discover that the accumulator for the mma(f32f8f8f32) instruction on the Ada architecture is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. Specifically, for mma(f32f8f8f32) instruction $C=AB+D$, where $A,B$ are tensors in FP8 data type and $C,D$ are tensors with FP32 data type, we initialize the $A,B$ to zero and vary $D$ to test the data type of the accumulator. As shown in Table 4, when $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ matches the result of the mma(f16f16f32) instruction. However, when $D$ is initialized with more than 13 mantissa bits, the error of $C$ corresponds to the difference between the results of mma(f32f16f16f32) and mma(f32f8f8f32). Consequently, the matrix multiplication of $\widetilde{P}V$, quantized to FP8, incurs a certain degree of accuracy loss compared to using an FP32 accumulator.
To mitigate this accuracy loss as much as possible, we propose to use an FP32 buffer to accumulate the values of $\widetilde{P}_{i}^{j}V_{j}$ in FP22. Specifically, as shown in Algorithm 1, after each Matmul between the tiled $\widetilde{P}$ and $V$, we store the Matmul result in the FP32 buffer. This approach prevents the FP22 accumulator from accumulating errors across the entire sequence, confining errors to the tiled range instead.
Table 5: Average accuracy using different data types of $(\widetilde{P},V)$ across all layers of a CogvideoX model, where $(Q,K)$ are smoothed.
INT4 INT8 77.05% 0.5618 0.5044
E5M2 99.20% 0.0905 0.0903
E4M3 99.44% 0.0683 0.0347
FP16 99.45% 0.0649 0.0335
@@ 3.4 FP32 MMA Buffer for FP22 Accumulator
Table 7: Two kernel implementations of SageAttention2.
      Kernel       $\psi_{Q}(Q)$, $\psi_{K}(K)$       $\psi_{P}(\widetilde{P})$, $\psi_{V}(V)$
      SageAttn2-4b       INT4, per-thread       FP8, per-block;  FP8, per-channel
      SageAttn2-8b       INT8, per-thread       FP8, per-block;  FP8, per-channel
3.5 Quantization for Q, K, P, V
Quantization for $Q,K$. We propose $\psi_{Q}(Q)$ and $\psi_{K}(K)$ in granularity of per-thread. This is first because per-channel quantization is not feasible since the scale factors of the inner axis of $QK^{\top}$ cannot used to do dequantization (Xiao et al., 2023a). Second, as shown in Table 3 and Table 3, we compare the average and worst-case accuracies of INT4 quantization at per-token, per-thread, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX. Results indicate that the accuracy of per-thread quantization is very close to per-token and outperforms a lot more than per-block and per-tensor. Furthermore, per-thread quantization incurs less dequantization overhead than per-token, as discussed in Section 3.3.
Quantization for $\widetilde{P},V$. We choose FP8, specifically the E4M3 data type, for $\psi_{P}(\widetilde{P})$ and $\psi_{V}(V)$ for two reasons. First, most GPUs have tensor cores that support FP8 Matmul operations, which are twice as fast as those using FP16. Second, Table 6 and Table 6 show the average and worst accuracy of different data types used for $\widetilde{P},V$ using real $Q,K,V$ across all layers of CogvideoX. We can see that the accuracy of using E4M3 is very close to that of using FP16 and superior to that of E5M2 and INT8. We propose to use $\psi_{P}(\widetilde{P})$ in per-block and $\psi_{V}(V)$ in per-channel for three reasons. First, per-channel quantization for $\widetilde{P}$ and per-token quantization for $V$ are not viable because dequantization requires scale factors of outer axis. Second, $\widetilde{P}=\mathrm{exp}(S_{i}-\mathrm{rowmax}(S_{i}))$, where $S_{i}$ is the Matmul result of a block of $Q$ and $K^{\top}$, the max value in each row of $\widetilde{P}$ is 1. Hence, we can assign a single static scale $s=\frac{1}{448}$ to a block $\widetilde{P}$. Third, per-channel quantization can address the channel-wised outlier of $V$.
3.6 An Optional Technique: Smooth V
We also figure out another way to mitigate the accuracy loss due to the FP22 accumulator when $V$ possesses channel-wise biases:
$\gamma(V)=V-\mathrm{mean}(V),$ $~{}~{}~{}~{}\mathrm{mean}(V)=\overrightarrow{V_{m}}$
$\phi_{V}(V)=$ $\psi_{V}\circ\gamma,~{}\overrightarrow{V_{m}}$ (7)
As shown in Figure 7, this strategy could enhance the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ consistently features channel-wise biases that are exclusively positive or negative, for instance, ranging between 8 and 9. Consequently, the values of $\widetilde{P}V$ could be quite large. However, the floating-point number representation range is not uniform—it is denser near zero. Therefore, by subtracting the mean along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in higher representational precision. We believe such an explanation and strategy could address lots of issues [^1] https://github.com/triton-lang/triton/issues/4476, https://github.com/triton-lang/triton/issues/5065 reported by the community. Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$. This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$. In other words, this method decomposes $V$ into two parts: $\overrightarrow{V_{m}}$ and $V$. For $V$, it centers the values of each column around zero, which leads to the dot product result between a row from the quantized $\widetilde{P}$ matrix and a column from the quantized $V$ matrix being closer to zero. This makes the representation of FP22 more accurate. Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.
Remark. This technique is optional and not employed in our main experiments. This is because its benefits are marginal, only significantly enhancing attention accuracy primarily when $V$ exhibits channel-wise outliers, which are absent in some models, such as $V$ in Llama3.1 (see Figure 4).
4 Experiment
4.1 Setup
Models. We validate the effectiveness of SageAttention2 across a diverse set of representative models from language, image, and video generation. Specifically, we conduct experiments on ten models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B), CogvideoX (1.5-5B) (Yang et al., 2024), HunyuanVideo (Weijie Kong & Jie Jiang, 2024), and Mochi (Team, 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) and Stable-Diffusion3.5 (turbo) (Stability AI, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
Datasets. Text-to-text models are evaluated on four zero-shot tasks: WikiText (Merity et al., 2022) to assess the model’s prediction confidence, LAMBADA (Paperno et al., 2016) evaluate contextual understanding, MMLU (Hendrycks et al., 2020) for measuring knowledge across various subjects, and Longbench (Bai et al., 2023) for evaluating bilingual, multitask, and comprehensive assessment of long context understanding capabilities. Text-to-video models are evaluated using the open-sora (Zheng et al., 2024) prompt sets. Text-to-image models are assessed on MJHQ-30K (Li et al., 2024). TIMM is evaluated on on three image datasets: ImageNet (Deng et al., 2009), ImageNet-Sketch (Sketch) (Wang et al., 2019), and ImageNet-Rendition (ImageNet-r) (Hendrycks et al., 2021).
Metrics. For text-to-text models, we use perplexity (ppl.) (Jelinek et al., 1977) for WikiText, Accuracy (Acc.) for LAMBADA and MMLU, and Longbench score (Bai et al., 2023). For text-to-video models, following Zhao et al. (2024), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024) to measure the text-video alignment; (VQA-a) and (VQA-t) to assess the video aesthetic and technical quality, respectively; and Flow-score (FScore) for temporal consistency (Wu et al., 2023). For text-to-image models, generated images are compared with the images in MJHQ-30K dataset in three aspects: FID (Heusel et al., 2017) and sFID (Salimans et al., 2016) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2024) for human preference. For TIMM, we use classification accuracy.
Implemetation details. We implement our attention kernels using CUDA and conduct experiments on Ubuntu 22.04 servers with RTX4090 and L20 GPUs.
Baselines. (1) SmoothAttn. Following Qserve (Lin et al., 2024), we apply smooth quant for matrixes $Q,K$ with smoothing factor $\alpha=0.5$. (2) HadmdAttn. Following Quarot (Ashkboos et al., 2024), we apply random Hadamard transformation for matrixes $Q,K$ before INT4 quantization. (3) SageAttention (Zhang et al., 2024) uses smoothing $K$, INT8 per-block quantization for $Q,K$, and FP16 for $\widetilde{P},V$.
Table 8: Average accuracy across all layers of CogvideoX using different smooth methods.
Method CosSim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
None 80.04% 0.3906 0.2223
HadmdAttn 79.77% 0.3782 0.2180
SmoothAttn 90.21% 0.3383 0.1952
Smooth K (Ours) 98.07% 0.1493 0.0743
Smooth Q (Ours) 98.30% 0.1250 0.0712
SageAttn2-4b 99.46% 0.0648 0.0334

### v4

@@ 4.1 Setup
Table 7: Average accuracy using different data types of $(\widetilde{P},V)$ across all layers of CogvideoX, where $(Q,K)$ are smoothed.
INT4 INT8 77.05% 0.5618 0.5044
E5M2 99.20% 0.0905 0.0903
E4M3 99.44% 0.0683 0.0347
FP16 99.45% 0.0649 0.0335
Model GPU Original SageAttn 2-8b SageAttn 2-4b
CogvideoX (2B) RTX4090 86 s 54 s 52 s
CogvideoX (1.5-5B) RTX4090 1040 s 577 s 555 s
HunyuanVideo L20 2221 s 1486 s 1435 s
Mochi L20 2336 s 1316 s 1190 s
Llama3.1 (48K token) RTX4090 9.2 s 5.7 s 5.6 s
Llama3.1 (100K token) L20 39.9 s 25.4 s 23.2 s
4.2 Speed and Accuracy of Kernels
Kernel Speed. We compare the Speed of SageAttention2 against baselines using configurations with headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017). Specifically, Fig. 5 shows the speed across varying sequence lengths on RTX4090, indicating that SageAttn2-4b and SageAttn2-8b are approximately 3x and 2.7x faster than FlashAttention2, and about 4.5x and 4x faster than xformers, respectively. Fig. 91011121314, and 15 in Appendix A.2 show more kernel speed comparison on RTX4090, L20, H20, H100 GPUs.
Accuracy. Table 4 and 17 show the average accuracy of different methods with INT4 $Q,K$ and FP8 $P,V$ across all layers of CogvideoX. The results indicate the accuracy of SageAttn2-4b is superior to other baselines.
4.3 End-to-end Performance
Metrics loss. We assessed the end-to-end metrics of various models using SageAttention2 compared to baselines. Detailed evaluation results are presented in Table 2. The results indicate that SageAttn2-4b outperforms all baselines and maintains most of the end-to-end accuracy across all models. Additionally, SageAttn2-8b incurs almost no metrics loss across various models. More experiment results on other models are shown in Appendix A.7.
Visible image and video examples. Fig.6 and Fig. 8 in Appendix A.1 show some visible comparision examples from HunyuanVideo, Mochi and CogvideoX (1.5-5B). We can observe that SageAttn2-8b does not introduce any visible differences compared to full-precision attention, whereas SageAttn2-4b has minor differences but is much better than the baselines.
4.4 Ablation Study
Overhead of techniques we proposed. As shown in Table 18, the overhead on kernel speed of per-thread quantization, smoothing Q, and two-level accumulation are 0.35%, 3.7%, and 0% compared to the attention kernel.
Benefit of smoothing V. The experiment showing the benefit of smoothing $V$ is shown in Appendix. A.4.
5 Conclusion
We introduce SageAttention2, an efficient and accurate quantized attention. First, we propose to quantize matrixes $(Q,K)$ in a thread-level granularity and $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth $Q$, enhancing the accuracy of $QK^{\top}$. Third, we propose a two-level accumulation strategy to enhance the accuracy of FP8 $\widetilde{P}V$. SageAttention2 is faster than FlashAttention2 and xformers by approximately 3x and 4.5x, respectively. Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, but offers significantly higher accuracy. Extensive testing confirms that our approach maintains end-to-end metrics across language, image, and video generation models.
Appendix A Appendix
A.1 Visible Comparison Exmaples
Figure 8: A comparison example using SageAttn2-8b and FlashAttention3 on CogvideoX-1.5, Mochi and HunyuanVideo.
A.2 Additional Kernel Speed Comparison
Fig. 9,  10,  11,  12,  13,  14, and  15 compare the speed of SageAttention2 against baselines using configurations with headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017), on RTX4090, L20, H100, and H20 GPUs.
Table 9 summarizes the performance gain of different attention methods against baselines on various modern GPUs.
Figure 12: Speed comparison between SageAttention2 and baselines (H100, headdim=64).
Figure 13: Speed comparison between SageAttention2 and baselines (H100, headdim=128).
Figure 14: Speed comparison between SageAttention2 and baselines (H20, headdim=64).
Figure 15: Speed comparison between SageAttention2 and baselines (H20, headdim=128).
Table 9: Speedup of different attention methods on various GPUs.
Method 3090 4090 A100 L40 L20 H100 H20
FlashAttention2 1.00 1.00 1.00 1.00 1.00 1.00 1.00
FlashAttention3 1.37 1.57
FlashAttention3 (fp8) 2.63 3.06
SageAttention1 1.97 1.96 1.37 1.45 1.24 1.53 1.52
SageAttention2 2.93 2.60 2.46 2.61 3.12
Smooth V Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
98.25% 0.1980 0.2387
99.75% 0.0406 0.0773
A.3 Smoothing V
As shown in Fig. 16, this strategy could enhance the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ in some models consistently features channel-wise biases that are exclusively positive or negative, for instance, ranging between 8 and 9 in CogvideoX. Consequently, the values of $\widetilde{P}V$ could be quite large. However, the floating-point number representation range is not uniform—it is denser near zero. Therefore, by subtracting the mean $\overrightarrow{V_{m}}$ along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in higher representational precision (see Fig. 16 for a visual demonstration). Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$. This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$. In other words, this method decomposes $V$ into two parts: $\overrightarrow{V_{m}}$ and $V$. For $V$, it centers the values of each column around zero, which leads to the dot product result between a row from the quantized $\widetilde{P}$ matrix and a column from the quantized $V$ matrix being closer to zero. This makes the representation of FP22 more accurate. Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.
A.4 Experiment of Smoothing V
Table 10 shows the attention accuracy on real tensors sampled from CogvideoX with and without smoothing $V$. It demonstrates that smoothing V could improve the accuracy of SageAttention2 when quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
A.5 Per-Thread Quantization Formulation
Figure 17: Memory layout of INT4/INT8 tensor core for accumulator matrix $C$ and $D$ in $D=A*B+C$ among 32 threads (T0 $\sim$ T31) in a warp. $C$ and $D$ is of shape 16x8. Each thread only holds 4 out of the 128 elements.
To further clarify the per-thread quantization, we first introduce the INT4 MMA instruction of Tensor Core, and then give the formulation of per-thread quantization.
Tensor cores, first introduced in NVIDIA’s Volta architecture, are specialized units designed for efficient matrix-multiply-and-accumulate (MMA) operations. Their usage significantly enhances computational efficiency and performance in AI and high-performance computing (HPC) workloads. Tensor cores compute small tiles of MMA operations, specifically $D=A*B+C$ on a warp (32 contiguous threads) basis. Each thread in the warp holds a fragment of input matrices and will get a fragment of output matrix as a computation result. The INT4 mma.m16n8k64 tensor core operation computes the product of a $16\times 64$ INT4 matrix $A$ and a $64\times 8$ INT4 matrix $B$, both stored in registers. It accumulates the result into a $16\times 8$ INT32 matrix $C$, also stored in registers, and returns the final product matrix $D$, which has the same shape ($16\times 8$), data type (INT32), and storage location (registers). Each thread holds only $\frac{1}{32}$ of the input and output data. Fig. 17 extracted from the PTX document (NVIDIA, ) shows the memory layout of matrix $C$ and $D$ among 32 threads in a warp. Each thread only holds 4 out of the 128 result elements.
$i_{\delta q}=$ $\lfloor(n*8*c_{w}/b_{q})\rfloor$
$q_{i}[i_{\delta q}]=$$\{8\times(n\%8)+\lfloor(n*\frac{c_{w}}{b_{q}})\rfloor*\frac{b_{q}}{c_{w}}\},~{}n\in[0,N]$
$\hat{Q}[q_{i}[i_{\delta q}]]=$ $\left\lceil\frac{Q[q_{i}[i_{\delta q}]]}{\delta_{Q}[i_{\delta q}]}\right\rfloor$
$i_{\delta k}=$ $\lfloor(n*4/b_{k})\rfloor$ (3)
$k_{n}[i_{\delta k}]=$$\{8\times(n\%8)+\lfloor n/b_{k}\rfloor*b_{k}\}\cup$
$\{8\times(n\%8)+1+\lfloor n/b_{k}\rfloor*b_{k}\},~{}~{}n\in[0,N]$
$\hat{K}[k_{n}[i_{\delta k}]]=$ $\left\lceil\frac{K[k_{n}[i_{\delta k}]]}{\delta_{K}[i_{\delta k}]}\right\rfloor$
In Eq. A.5, $c_{w}$ is the count of GPU Warps. $b_{q}$ and $b_{k}$ are the block size of $Q,K$. $n$ is the token index of $Q,K$.
By ensuring results held by each thread share a common dequantization scale (belong to the same quantization group), we can avoid the overhead associated with per-token quantization. Leveraging this observation, we design per-thread quantization, as shown in Fig. 4. For typical block size of $b_{q}=128$, $b_{k}=64$ and warp number $c_{w}=4$ (as used in FlashAttention2), each warp processes a tile of 32 query tokens and 64 key tokens. Query tokens $i,8+i,16+i,24+i$ ($i=0,1,\cdots,7$) can be made into one quantization group and key tokens $j,1+j,8+j,9+j,\cdots,56+j,57+j$ ($j=0,1,2,3$) can be made into one quantization group. This design aligns with the memory layout of output matrix $D$ of tensor core shown in Fig. 17, ensuring that each thread only needs one $Q$ scale and one $K$ scale for dequantization.
As a result, this approach creates 32 quantization groups for $Q$ (8 for each of the 4 warps) and 4 quantization groups for $K$ in a 128x64 block, providing 32×32$\times$32 × and 4×4$\times$4 × finer granularity compared to per-block quantization for query tokens and key tokens, respectively. Table 6 and Table 15 show the accuracy gains by using per-thread quantization. Per-thread quantization achieves accuracy that closely matches per-token quantization, without introducing any kernel speed degradation (see Fig. 18).
A.6 Datasets and Metrics in Experiments
Datasets. Text-to-text models are evaluated on four zero-shot tasks: WikiText (Merity et al., 2022) to assess the model’s prediction confidence, LAMBADA (Paperno et al., 2016) evaluate contextual understanding, MMLU (Hendrycks et al., 2020) for measuring knowledge across various subjects, and Longbench (Bai et al., 2024) for comprehensive assessment of long context understanding capabilities. Text-to-video models are evaluated using the open-sora (Zheng et al., 2024) prompt sets. Text-to-image models are assessed on MJHQ-30K (Li et al., 2024). TIMM is evaluated on on three image datasets: ImageNet (Deng et al., 2009), ImageNet-Sketch (Sketch) (Wang et al., 2019), and ImageNet-Rendition (ImageNet-r) (Hendrycks et al., 2021).
End-to-end metrics. For text-to-text models, we use perplexity (ppl.) (Jelinek et al., 1977) for WikiText, Accuracy (Acc.) for LAMBADA and MMLU, and Longbench score (Bai et al., 2024). For text-to-video models, following Zhao et al. (2025), we evaluate the quality of generated videos on five metrics: CLIPSIM and CLIP-Temp (CLIP-T) (Liu et al., 2024) to measure the text-video alignment; (VQA-a) and (VQA-t) to assess the video aesthetic and technical quality, respectively; and Flow-score (FScore) for temporal consistency (Wu et al., 2023). For text-to-image models, generated images are compared with the images in MJHQ-30K dataset in three aspects: FID (Heusel et al., 2017) and sFID (Salimans et al., 2016) for fidelity evaluation, Clipscore (CLIP) (Hessel et al., 2021) for text-image alignment, and ImageReward (IR) (Xu et al., 2023) for human preference. For TIMM, we use classification accuracy.
Table 11: End-to-end metrics on Llama2 (7B).
Model Attention WikiText (Ppl.) $\downarrow$ Lambda (Acc.) $\uparrow$ MMLU (Acc.) $\uparrow$
Llama2 Full-Precision 5.823 0.886 0.439
HadmdAttn 6.771 0.860 0.360
SmoothAttn 6.717 0.867 0.392
SageAttention 5.824 0.887 0.439
SageAttn2-4b 5.912 0.881 0.428
SageAttn2-8b 5.828 0.886 0.438
Table 12: End-to-end metrics on CogvideoX (2B).
CogvideoX (2B) Full-Precision 0.1836 0.9975 77.605 75.360 3.006
HadmdAttn 0.1742 0.9877 29.780 23.985 0.499
SmoothAttn 0.1741 0.9870 41.703 47.043 0.624
SageAttention 0.1833 0.9976 76.997 71.360 2.988
SageAttn2-4b 0.1821 0.9973 77.368 74.906 2.603
SageAttn2-8b 0.1829 0.9977 76.532 74.281 2.941
Table 13: End-to-end metrics on an image classification model.
TIMM Full-Precision 84.79% 45.32% 59.55%
HadmdAttn 84.50% 44.89% 58.80%
SmoothAttn 84.40% 44.68% 58.73%
SageAttention 84.74% 45.38% 59.95%
SageAttn2-4b 86.67% 45.24% 59.29%
SageAttn2-8b 84.79% 45.39% 59.57%
Table 14: Comparison with FlashAttention3(fp8) on Llama-3-262k (8B) on InfiniBench (Zhang et al., 2024) (H100 GPU).
Attention Eng.Sum Eng.QA Eng.MC Code.Debug Math.Find Retr.PassKey Retr.Num Retr.KV Avg.
Full-Precision 18.03 12.5 64.19 24.37 18.29 100.0 100.0 7.0 43.05
FlashAttn3-fp8 19.03 11.73 55.90 22.59 22.57 100.0 100.0 0.4 41.53
SageAttention2 18.17 12.46 64.19 25.63 17.43 100.0 100.0 6.6 43.06
(a) Full Precision
(b) FlashAttn3-fp8
(c) SageAttention2
Figure 18: Needle In A Haystack results on Llama-3-262k (8B).